OpenStudy (haleyelizabeth2017):

Simplify the rational expression. State any restrictions on the variable.

OpenStudy (haleyelizabeth2017):

\[\frac {n^4-11n^2+30} {n^4-7n^2+10}\] I have it simplified, I just need help finding restrictions ;) \(\Large\frac {(n^2-6)(n^2-5)} {n^2-2)(n^2-5)}\) = \(\Large\frac {n^2-6} {n^2-2}\)

jimthompson5910 (jim_thompson5910):

Look at the original expression. Specifically at the denominator when it is factored.

jimthompson5910 (jim_thompson5910):

When does it equal 0?

OpenStudy (haleyelizabeth2017):

One second ;)

OpenStudy (haleyelizabeth2017):

Thanks! :D

jimthompson5910 (jim_thompson5910):

what 4 values of n were you able to find?